# Part I: General framework, application to Hartree-Fock and DFT

### v1

@@ 5.2 Multiple critical points within the UHF theory for H4
Due to the nonlinearities induced by the UHF approximation, there can be multiple critical points (in particular, index-1 saddle points) over the UHF electronic landscape. In this subsection, we consider the H4 molecule in a rectangular configuration with side lengths of 1 Å and 1.1 Å, at the UHF/STO-3G 57 level of theory. The Hamiltonian is parametrized by the coupling parameter $\lambda$, as in Section 4.1. For each $\lambda\in[0,2]$ with a spacing of $0.05$, we find the index-1 saddle points by running the saddle search algorithm from 50 different randomly generated initial points over the Grassmannian.
@@
Fig. 6 shows the energies of the index-1 saddle points which appear in the range $\lambda\in[0,2]$.
@@
Figure 6: Dependence on the $\lambda$-parameter of the energy of the index-1 saddle points, calculated on H4 in a rectangular configuration with sides of 1 Å and 1.1 Å. The UHF/STO-3G level of theory has been used. Calculations have been made in steps of $\Delta\lambda=0.2$.
@@
We see that in this range, at most three critical points are found. In the range $\lambda\in[0,0.35]$, only one index-1 saddle point is found, in $\lambda\in[0.40,0.60]$ two index-1 saddle points are found, and from there with increasing $\lambda$, three index-1 saddle points are observed. At $\lambda\approx 1.68$ the lines corresponding to the second and third saddle points are crossing, showing that the ordering in the energies of the states might not be preserved.
@@
It is known that not all index-1 saddle points provide suitable approximations for excited states. An interpretation of the critical points can be gained through projection onto the FCI states, as these form a complete basis. Our calculations consider the case of $S_{z}=0$, therefore fixing $N_{\alpha}=N_{\beta}=2$. In the present case of a minimal basis set, a total number of 36 FCI states can be built. The projection of the wave functions corresponding to the critical points can be written as
|$\ket{\Psi^{\mathrm{CP-HF}}(\lambda)}=\sum_{I}c_{I}(\lambda)\ket{\Psi_{I}^{\mathrm{FCI}}(\lambda)}.$
@@
The magnitudes of the $c_{I}(\lambda)$ coefficients indicate the extent to which the critical points correspond to the $I$-th FCI state. For the sake of clarity, we will include only the eight lowest-lying FCI states, as the contributions of the higher-energy states are relatively small. We include the evolution of the considered FCI excited energies with the coupling parameter in Fig. 7(a).
@@
Figure 7: (a) Dependence on the $\lambda$-parameter of the FCI energies, and (b-d) dependence on the $\lambda$-parameter of the overlaps of the UHF critical points with the FCI states, calculated on H4 in a rectangular configuration with sides of 1 Å and 1.1 Å. Calculations have been made in steps of $\Delta\lambda=0.05$.
@@
We label these states as $\Psi_{0}^{\mathrm{FCI}}(\lambda)$, $\Psi_{1}^{\mathrm{FCI}}(\lambda)$, $\ldots$, $\Psi_{7}^{\mathrm{FCI}}(\lambda)$ in accordance to their energy ordering in the non-interacting case. At $\lambda=0$, we observe that the pairs ($\Psi_{1}^{\mathrm{FCI}}(0)$, $\Psi_{2}^{\mathrm{FCI}}(0)$), ($\Psi_{4}^{\mathrm{FCI}}(0)$, $\Psi_{5}^{\mathrm{FCI}}(0)$), and ($\Psi_{6}^{\mathrm{FCI}}(0)$, $\Psi_{7}^{\mathrm{FCI}}(0)$) are degenerate. These degeneracies can be understood by considering the occupation of the molecular orbitals. In the non-interacting case, the lowest eight FCI states are either singlets or triplets and are therefore either single Slater determinants or linear combinations of two Slater determinants:
|$\ket{\Psi_{0}^{\rm FCI}(0)}$ =|$=\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{2}^{\alpha}\phi_{2}^{\beta}}$
(|$(\ket{\Psi_{1}^{\rm FCI}(0)},\ket{\Psi_{2}^{\rm FCI}(0)})$ $=\frac{1}{\sqrt{2}}(\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{2}^{\alpha}\phi_{3}^{\beta}}\pm\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{2}^{\beta}\phi_{3}^{\alpha}})$
|$\ket{\Psi_{3}^{\rm FCI}(0)}$ =|$=\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{3}^{\alpha}\phi_{3}^{\beta}}$
(|$(\ket{\Psi_{4}^{\rm FCI}},\ket{\Psi_{5}^{\rm FCI}(0)})$ $=\frac{1}{\sqrt{2}}(\ket{\phi_{1}^{\alpha}\phi_{2}^{\beta}\phi_{2}^{\alpha}\phi_{3}^{\beta}}\pm\ket{\phi_{1}^{\beta}\phi_{2}^{\beta}\phi_{2}^{\alpha}\phi_{3}^{\alpha}})$
(|$(\ket{\Psi_{6}^{\rm FCI}(0)},\ket{\Psi_{7}^{\rm FCI}(0)})$ $=\frac{1}{\sqrt{2}}(\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{2}^{\alpha}\phi_{4}^{\beta}}\pm\ket{\phi_{1}^{\alpha}\phi_{1}^{\beta}\phi_{2}^{\beta}\phi_{4}^{\alpha}})$
@@
The degeneracies arise from the fact that at $\lambda=0$, the singlet state and the triplet state of $S_{z}=0$ have the same energy, as discussed in Section 4.1.
@@
For $\lambda>0$, it can be observed that the degeneracy is lifted, as the singlet and triplet states no longer give the same energy. With increasing interaction parameter $\lambda$, the FCI states are observed to cross. In particular, $\Psi_{2}^{\mathrm{FCI}}$ (green line in Fig. 7(a)) crosses consecutively with the states $\Psi_{3}^{\mathrm{FCI}}$, $\Psi_{4}^{\mathrm{FCI}}$, $\Psi_{5}^{\mathrm{FCI}}$, $\Psi_{6}^{\mathrm{FCI}}$ (respectively, red, purple, brown, and pink lines in Fig. 7(a)) at $\lambda\approx 0.77$, 1.33, 1.49, and 1.94 respectively, as indicated by the dot, square, star, and diamond markers in the figure.
In Figs. 7(b)-7(d), the $c_{I}$ coefficients corresponding to the three index-1 saddle points are plotted. In Fig. 7(b), the analysis is performed for the lowest-lying critical point (CP1, blue line in Fig. 6). This saddle point appears only from $\lambda\approx 0.4$ and shows to be dominated by the FCI ground state. Therefore, this critical point can be interpreted as a spurious saddle point arising from the use of an approximate wave function method (UHF), which introduces nonlinearities. The existence of spurious critical points in nonlinear approximation models has been reported many times since the early days of quantum chemistry, and the study of this phenomenon is still an active field of research (cf. e.g., the recent work by Marie and Burton in Ref. 7 on CASSCF). Fig. 7(c) shows the same analysis for the second critical point (CP2, orange line in Fig. 6). This state remains a combination of $\Psi_{1}^{\rm FCI}$ and $\Psi_{2}^{\rm FCI}$. In particular, at $\lambda=0$, we have |$\ket{\Psi_{2}^{\mathrm{HF,CP}}(0)}=\frac{1}{\sqrt{2}}(\ket{\Psi_{1}^{\mathrm{FCI}}(0)}+\ket{\Psi_{2}^{\mathrm{FCI}}(0)})$ at $\lambda=0$. As the two states $\Psi_{1}^{\rm FCI}$ and $\Psi_{2}^{\rm FCI}$ are degenerate at $\lambda=0$ but belong to different spin-symmetry sectors, this critical point is a symmetry-broken solution, formed by a linear combination of the singlet first excited state $\Psi_{1}^{\rm FCI}$ and the triplet ground state $\Psi_{2}^{\rm FCI}$.
In Fig. 7(d), the overlap of the third critical point (CP3, green line in Fig. 6) is shown. It is observed that this critical point, which arises only from $\lambda=0.60$, has a large overlap with the state $\Psi_{3}^{\mathrm{FCI}}$. In particular, for $\lambda\in[0.60,1]$ the overlap is very close to one, but then the saddle point becomes a linear combination of states $\Psi_{3}^{\mathrm{FCI}}$ and $\Psi_{5}^{\mathrm{FCI}}$.

### v2

@@ 5.2 Multiple critical points within the UHF theory for H4
Due to the nonlinearities induced by the UHF approximation, there can be multiple critical points (in particular, index-1 saddle points) over the UHF electronic landscape. In this subsection, we consider the H4 molecule in a rectangular configuration with side lengths of 1 Å and 1.1 Å, at the UHF/STO-3G[71] level of theory. The Hamiltonian is parametrized by the coupling parameter $\eta$, as in Section 4.1. For each $\eta\in[0,2]$ with a spacing of $0.05$, we find the index-1 saddle points by running the saddle search algorithm from 50 different randomly generated initial points over the Grassmannian.
@@
Fig. 7 shows the energies of the index-1 saddle points which appear in the range $\eta\in[0,2]$.
@@
Figure 7: Dependence on the $\eta$-parameter of the energy of the index-1 saddle points, calculated on H4 in a rectangular configuration with sides of 1 Å and 1.1 Å. The UHF/STO-3G level of theory has been used. Calculations have been made in steps of $\Delta\eta=0.2$.
@@
We see that in this range, at most three critical points are found. In the range $\eta\in[0,0.35]$, only one index-1 saddle point is found, in $\eta\in[0.40,0.60]$ two index-1 saddle points are found, and from there with increasing $\eta$, three index-1 saddle points are observed. At $\eta\approx 1.68$ the lines corresponding to the second and third saddle points are crossing, showing that the ordering in the energies of the states might not be preserved.
@@
It is known that not all index-1 saddle points provide suitable approximations for excited states. An interpretation of the critical points can be gained through projection onto the FCI states, as these form a complete basis. Our calculations consider the case of $S_{z}=0$, therefore fixing $N_{\alpha}=N_{\beta}=2$. In the present case of a minimal basis set, a total number of 36 FCI states can be built in the $S_{z}=0$ sector of the FCI space. The projection of the wave functions corresponding to the critical points can be written as
[…]
@@
The magnitudes of the $c_{I}(\eta)$ coefficients indicate the extent to which the critical points correspond to the $I$-th FCI state. For the sake of clarity, we will include only the eight lowest-lying FCI states, as the contributions of the higher-energy states are relatively small. We include the evolution of the considered FCI excited energies with the coupling parameter in Fig. 8(a).
@@
Figure 8: (a) Dependence on the $\eta$-parameter of the FCI energies, and (b-d) dependence on the $\eta$-parameter of the overlaps of the UHF critical points with the FCI states, calculated on H4 in a rectangular configuration with sides of 1 Å and 1.1 Å. Calculations have been made in steps of $\Delta\eta=0.05$.
@@
We label these states as $\Psi_{0}^{\mathrm{FCI}}(\eta)$, $\Psi_{1}^{\mathrm{FCI}}(\eta)$, $\ldots$, $\Psi_{7}^{\mathrm{FCI}}(\eta)$ in accordance to their energy ordering in the non-interacting case. At $\eta=0$, we observe that the pairs ($\Psi_{1}^{\mathrm{FCI}}(0)$, $\Psi_{2}^{\mathrm{FCI}}(0)$), ($\Psi_{4}^{\mathrm{FCI}}(0)$, $\Psi_{5}^{\mathrm{FCI}}(0)$), and ($\Psi_{6}^{\mathrm{FCI}}(0)$, $\Psi_{7}^{\mathrm{FCI}}(0)$) are degenerate. These degeneracies can be understood by considering the occupation of the molecular orbitals. In the non-interacting case, the lowest eight FCI states are either singlets or triplets and are therefore either single Slater determinants or linear combinations of two Slater determinants:
[…]
@@
The degeneracies arise from the fact that at $\eta=0$, the singlet state and the triplet state of $S_{z}=0$ have the same energy, as discussed in Section 4.1.
@@
For $\eta>0$, it can be observed that the degeneracy is lifted, as the singlet and triplet states no longer give the same energy. With increasing interaction parameter $\eta$, the FCI states are observed to cross. In particular, $\Psi_{2}^{\mathrm{FCI}}$ (green line in Fig. 8(a)) crosses consecutively with the states $\Psi_{3}^{\mathrm{FCI}}$, $\Psi_{4}^{\mathrm{FCI}}$, $\Psi_{5}^{\mathrm{FCI}}$, $\Psi_{6}^{\mathrm{FCI}}$ (red, purple, brown, and pink lines in Fig. 8(a)) at $\eta\approx 0.77$, 1.33, 1.49, and 1.94 respectively, as indicated by the dot, square, star, and diamond markers in the figure.
[…]
@@
6 Conclusions
In this work, a unified and compact framework, based on the concept of Kähler manifolds, has been presented for the description of excited states in electronic structure theory, covering two different approaches, critical point search (CP) and linear response theory (LR). In particular, we describe the framework in a pedagogical way over the Grassmann manifolds, which amounts to be a more straightforward and systematic formulation of LR than Casida’s derivations when dealing with the mean-field models.
It was shown that CP and LR coincide for the FCI level of theory, but give different formulations for nonlinear approximations of the wave function. For the mean-field models such as HF and DFT, the CP approach amounts to finding different solutions of the HF or the KS equations, while the LR approach boils down to solving Casida’s equations, commonly used in LR-TDHF and LR-TDDFT. We have pointed out that TDHF and TDDFT can only target single excitations. On the contrary, the CP approach can target higher excitations and allows for orbital relaxation, at the cost of introducing additional or even non-physical candidates for excited states.
To analyze the differences between the CP and LR approaches, we have explored the weakly-interacting regime, based at the non-interacting limit, where the energy functional becomes linear and the LR and CP approaches are equivalent to compute the lowest excitation energies regardless of the level of theory in use. In particular, we exploit the perturbation theory to find the analytic first-order derivatives of excitation energies with respect to an additional coupling parameter, controlling the amount of interaction. For example in the UHF level of theory, we find that the CP approach yields a derivative that differs from the one in the FCI level, while the LR approach gives the correct first-order expression for the weakly-interacting case.
Preliminary numerical simulations have been performed to show the comparison of the CP and LR approaches in the UHF level of theory as well as the FCI, on the chosen systems H2, H4, and H2O, focusing on the first excited state. For the CP approach, the search is restricted to index-1 saddle points. The simulations have confirmed the analytic expressions for the first-order excitation energies at the non-interacting regime. Furthermore, we have analyzed the case of H4, where multiple index-1 saddle points could be identified. The chemical interpretation of these points has been achieved through projection on the basis of the FCI states. In this analysis, it has been shown that some spurious index-1 saddle points arise from the introduction of nonlinearities in the linear Schrödinger equation. Only one state could be shown to correspond to an excited state, a symmetry-broken solution given by the combination of the first-excited singlet and triplet states. Therefore, some caution should be adopted when trying to assign physical meanings to the critical points of nonlinear approximation models.
A more accurate description of excited states can be achieved by post-HF methods. In particular, we aim at analyzing the CASSCF manifold through the presented theory based on Kähler manifolds. The formalization of CP and LR descriptions of CASSCF excited states will be the matter of Part II.